The perimeter of a yield sign in the shape of an isosceles triangle is 22 feet. If the shortest side is 2 feet less than the other two sides, find the length of the shortest side.
Answer by vksarvepalli(154)

let the length of the sides be a,a and b
given
perimeter a+b+a=22 => 2a+b=22
and b=a-2 substituting this in the above eq.
we get 2a+a-2=22
=> 3a=24
=> a=8
so length of the shortest side = b= 8-2 = 6
length of the shortest side = 6 feet